OpenStudy (anonymous):

Write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2,-2/3)

OpenStudy (anonymous):

Use the vertex form of the parabola: y = a(x-xV)^2 + yV y = a(x - 1/2)^2 - 2/3 plug in 1 of the roots 0 = a(0 - 1/2)^2 - 2/3 0 = a/4 - 2/3 a/4 = 2/3 a = 8/3 the equation of the parabola: y = 8/3(x - 1/2)^2 - 2/3 or y = 8/3*x^2 - 8/3*x

OpenStudy (anonymous):

medal me if helped

OpenStudy (kvng_sj):

welcome to open study!! but okay whats the options example A,B,C,D?